# Coordination Requires Simplification: Thermodynamic Bounds on Multi-Objective Compromise in Natural and Artificial Intelligence

### v1

@@ IV.4 Energy for Metastable Coordination
Figure 1: Thermodynamic Coordination Theory: Key Relationships. (a) Information scaling showing $\log_{10}$ of coordination protocol length $L(P)$ as a function of number of agents $N$ and objectives $d$. Cyan dashed line marks human working memory limit (100 bits), demonstrating how even modest multi-agent coordination exceeds human capacity. Contours show information growth in orders of magnitude. Parameters: mean agent model complexity $\bar{K}=10$ bits, conflict resolution precision $\varepsilon=0.1$. (b) Distribution of solutions across agent space. Each cell represents one agent and the color represents which solution(s) they can find and agree to implement. Findable solutions (green+yellow) form extended connected regions reaching into corners; accurate but “invisible” solutions (red+yellow) exist as isolated islands. Yellow regions indicate where both solution types are available. The green+yellow network forms a spanning cluster enabling system-wide coordination, while red+yellow regions remain disconnected. This collective marginal utility increase drives selection for findable solutions over more accurate one, connecting to topological constraints from Arrow’s impossibility theorem (see Supplementary Section I). (c) Coordination temperature states visualized through variance in agent models. Each circle represents one agent’s internal model state. Circular clusters from left to right show: ordered state with tightly aligned models (low $T_{\mathrm{co}}$), critical coordination state with moderate variance, and disordered state with maximal model divergence (high $T_{\mathrm{co}}$). (d) Minimum thermodynamic work required to cool system from initial temperature $T_{1}$ to final temperature $T_{2}(<T_{1})$. Bottom axis shows temperature ratio $T_{2}/T_{1}$; top axis shows absolute final temperature $T_{2}$ assuming $T_{1}=1$. Curve height represents work in bits. Critical temperature $T_{c,\mathrm{co}}\approx 0.128$ determines stability: systems with $T_{2}<T_{c,\mathrm{co}}$ (green region) can maintain coordination without additional energy input; systems with $T_{2}>T_{c,\mathrm{co}}$ (red region) require continuous resource expenditure for metastable coordination. Parameters: $N=50$, $\bar{K}=20$, $K_{0}=10$ bits.
V Discussion
@@
V.1 Empirical Validation
@@
Artificial Intelligence: The indefinite cycling observed in multi-objective gradient descent [e.g. 26] directly supports this framework, in particular the extension of Chichilnisky-Arrow’s Impossibility Theorem. While different “gradient surgery” methods have been proposed to manage conflicting objectives, and indeed do improve convergence, the problem remains algorithmically intractable. Increased computational capacity cannot escape these bounds as larger systems require coordination among more agents (annotators, users, developers) with more objectives, causing $N^{2}d^{2}$ to dominate any linear capacity increase, i.e. computational scaling laws do not scale [43]. Reinforcement learning from human feedback could be making the problem worse [44], and leading to alignment faking [45]. Large Language Models (LLM) provide the ideal conditions to test the TCT framework with comparatively simpler ways to measure the parameters needed ($N,\ d,\ K,\ \varepsilon,\ m_{i}$). Detailed analysis of these results and their implications for machine learning systems are topics for further work.
@@
Other Systems: Empirical signatures of TCT appear across different contexts and hierarchical levels. Human reasoning systematically favors coordination over accuracy [46], with more education paradoxically leading to worse identity protective cognition [47]. Arrow’s impossibility appears to extend to an individual’s own preferences, which can be reversed by simple changes in how choices are framed [48], suggesting the challenge of consistent preference aggregation arises from the brain’s own internal ’society’ of competing subsystems even before the multi-agent coordination problem begins [49]. This observation originally motivated this theoretical framework to explain the similar phenomena in LLMs where scaled systems converge to generic outputs [45, 44, 43] [^2] Citations to pre-prints are included due to the rapid pace of research in machine learning.. Cultural evolution selects for transmissible simplifications [50, 51]. Scientific paradigm appears to shift in episodic “Kuhn cycles” instead of development by accumulation [52]. We predict open source projects are expected to show incremental API complexity decreasing as a power-law: $K_{\rm{API}}\propto N^{-\alpha}$ where $\alpha\in(0,1)$, reflecting hierarchical coordination requirements (Equation III.1).
@@
Market bubbles show critical transitions [53] potentially predictable from trader density, an analysis subject to future work. The difference in the description length of model and communication complexity terms in Equation III.1 potentially explains the paradox noted in behavioral finance: why do similarly trained professionals make opposite trades on identical information [54]? High model overlap does not imply preference alignment.
@@
V.2 Scale Free Implications
@@ V.3 Limitations
All systems which successfully survive and propagate most likely discover workable ways around paying the full $N^{2}d^{2}$ coordination costs, such as using hierarchical coordination protocols (Equation III.1).
@@
V.4 Suggested System Design Strategies

### v3

@@ 4.4 Energy for Metastable Coordination
Figure 1: Thermodynamic Coordination Theory: Key Relationships. (a) Information scaling showing $\log_{10}$ of coordination protocol length $L(P)$ as a function of number of agents $N$ and objectives $d$. Cyan dashed line marks human working memory limit (100 bits), demonstrating how even modest multi-agent coordination exceeds human capacity. Contours show information growth in orders of magnitude. Parameters: mean agent model complexity $\bar{K}=10$ bits, conflict resolution precision $\varepsilon=0.1$. (b) Distribution of solutions across agent space. Each cell represents one agent and the color represents which solution(s) they can find and agree to implement. Findable solutions (green+yellow) form extended connected regions reaching into corners; accurate but “invisible” solutions (red+yellow) exist as isolated islands. Yellow regions indicate where both solution types are available. The green+yellow network forms a spanning cluster enabling system-wide coordination, while red+yellow regions remain disconnected. This collective marginal utility increase drives selection for findable solutions over more accurate one, connecting to topological constraints from Arrow’s impossibility theorem (see Supplementary Section 1). (c) Coordination temperature states visualized through variance in agent models. Each circle represents one agent’s internal model state. Circular clusters from left to right show: ordered state with tightly aligned models (low $T_{\mathrm{co}}$), critical coordination state with moderate variance, and disordered state with maximal model divergence (high $T_{\mathrm{co}}$). (d) Minimum thermodynamic work required to cool system from initial temperature $T_{1}$ to final temperature $T_{2}(<T_{1})$. Bottom axis shows temperature ratio $T_{2}/T_{1}$; top axis shows absolute final temperature $T_{2}$ assuming $T_{1}=1$. Curve height represents work in bits. Critical temperature $T_{c,\mathrm{co}}\approx 0.128$ determines stability: systems with $T_{2}<T_{c,\mathrm{co}}$ (green region) can maintain coordination without additional energy input; systems with $T_{2}>T_{c,\mathrm{co}}$ (red region) require continuous resource expenditure for metastable coordination. Parameters: $N=50$, $\bar{K}=20$, $K_{0}=10$ bits.
5 Discussion
@@
5.1 Cross-Domain Signatures
We identify signatures in existing observations as preliminary validation:
Artificial Intelligence: The indefinite cycling observed in multi-objective gradient descent (e.g. [26]) directly supports this framework, in particular the extension of Chichilnisky-Arrow’s Impossibility Theorem. Each loss function $L_{i}(\theta)$ defines a preference ordering over parameter space, and optimizing $d$ potentially conflicting objectives requires aggregating these orderings into a single gradient update direction, typically through weighted combination: $\nabla\theta=\sum\alpha_{i}\nabla L_{i}$. The selection of weights $\{\alpha_{i}\}$ is topologically equivalent to Chichilnisky’s aggregation function $F$: it must map $d$ preference orderings to a collective preference (update direction). The impossibility result applies directly as no continuous, anonymous, unanimous weighting scheme exists when objectives conflict. While different “gradient surgery” methods have been proposed to manage conflicting objectives and provide practical improvements, the problem remains algorithmically intractable. The system cannot converge because no consistent aggregation exists, only arbitrary compromises that change under perturbation (e.g. [43]).
@@ 5.1 Cross-Domain Signatures
Increased computational capacity cannot escape these bounds as larger systems require coordination among more agents (annotators, users, developers) with more objectives, causing $N^{2}d^{2}$ to dominate any linear capacity increase, i.e. computational scaling laws do not scale [44]. Reinforcement learning from human feedback could be making the problem worse [45], and leading to alignment faking [46]. Large Language Models (LLM) provide the ideal conditions to test the TCT framework with comparatively simpler ways to measure the parameters needed ($N,\ d,\ K,\ \varepsilon,\ m_{i}$). Detailed analysis of these results and their implications for machine learning systems are topics for further work.
@@
Other Systems: Qualitative signatures of TCT appear across different contexts and hierarchical levels. Human reasoning systematically favors coordination over accuracy [47], with more education paradoxically leading to worse identity protective cognition [48] to defend focal points. Arrow’s impossibility appears to extend to an individual’s own preferences, which can be reversed by simple changes in how choices are framed [49], suggesting the challenge of consistent preference aggregation arises from the brain’s own internal ’society’ of competing subsystems even before the multi-agent coordination problem begins [50]. This observation originally motivated this theoretical framework to explain the similar phenomena in LLMs where scaled systems converge to generic outputs [46, 45, 44] [^2] Citations to pre-prints presented at conferences are included due to the rapid pace of research in machine learning.. In these contexts, the focal points that emerge from coordination can be understood as the high-level priors that guide the predictive processing of the collective system.
Cultural evolution selects for transmissible simplifications [51, 52]. Scientific paradigm appears to shift in episodic “Kuhn cycles” instead of development by accumulation [53]. We predict open source projects are expected to show incremental API complexity decreasing as a power-law: $K_{\rm{API}}\propto N^{-\alpha}$ where $\alpha\in(0,1)$, reflecting hierarchical coordination requirements (Equation 9).
@@
Market bubbles show critical transitions [54] potentially predictable from trader density, an analysis subject to future work. The difference in the description length of model and communication complexity terms in Equation 3.1 potentially explains the paradox noted in behavioral finance: why do similarly trained professionals make opposite trades on identical information [55]? High model overlap does not imply preference alignment.
@@
5.2 Scale Free Implications
@@ 5.3 Limitations
All systems which successfully survive and propagate most likely discover workable ways around paying the full $N^{2}d^{2}$ coordination costs, such as using hierarchical coordination protocols (Equation 9).
@@
5.4 Suggested System Design Strategies
@@ 5.5 Conclusions
[…]
No new data was created or analyzed in this study.
[…]
Supplementary PDF includes the topological derivation of the Theorem 1, derives the minimum description length of communication protocols in hierarchical coordination, and details characterization of Level II chaos dynamics using the bistromathics (restaurant bill splitting) example.